OpenStudy (anonymous):

Solve for x. 3x = y + 4 2x + 9 - 2y = y

OpenStudy (anonymous):

does the second one really say \(2x+9-2y=y\) ? ?

OpenStudy (anonymous):

Yes. Yes it does XD

OpenStudy (anonymous):

if so, lets clean that up first by rewriting it as \(2x-3y=-9\)

OpenStudy (anonymous):

and if \(3x = y + 4\) then we can write \(y=3x-4\) and replace \(y\) in the second equation by \(3x-4\) with judicious use of parentheses to get \[2x-3(3x-4)=-9\] and solve that one for \(x\)

OpenStudy (anonymous):

you good from there or should we go ahead and solve the whole thing?

OpenStudy (anonymous):

I understand now! Thank you c: It's just the whole placing of the numbers was confusing!

OpenStudy (anonymous):

ok good if you want, post your answer and i can check it

OpenStudy (anonymous):

y=5 x=3??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you c:

OpenStudy (anonymous):

yw